The perimeter of Triangle ABC is 306 inches. Find the lengths of the sides.
My problem is that I am not sure what formula to use.
AB=_____ in BC=_______in AC=______in

Answer by checkley77(12844)

Without any additional information there are an infinite number of solutions to this probnlem.
a+b+c=306
102+102+102=306 is one solution.
103+101+102=306 is another solution.
101+101+104=306 is a third solution.
104+104+98=306 is a fourth solution.
And there are many many more.